How many lattice points are there in one unit cell of each of the following lattice? (i) Face-centred cubic (ii) Face-centred tetragonal (iii) Body-centred

Grade:8

1 Answers

Gaurav
9 years ago
Hello Student
(i) A face–centred cubic (fcc)unit cell contains atoms at all the corners and at the centre of all the faces of the cube.
So lattice point at the corner= 8
Lattice point at the center of all faces= 6
Total number of lattice point= 14
(ii) Face–centred tetragonal
lattice point at the corner= 8
Lattice point at the center of all faces= 6
Total number of lattice point = 14
(iii) Body–centred
In Body centered cubic there are 8 lattice point at the corner and 1 is at the center so total 8+1 =9 lattice point